# Federated Learning Privacy: Attacks, Defenses, Applications, and Policy Landscape — A Survey

### v1

@@ 7.3. Homomorphic Encryption
Clients can employ HE to encrypt their model updates within the FL paradigm before transmission. Consequently, the server exclusively receives encrypted updates and is precluded from accessing any raw model updates. Utilizing an HE protocol, the server aggregates only encrypted weights, thereby preserving privacy. Following aggregation, the updated global model is sent back to the clients while still being encrypted. Subsequently, clients decrypt the model using their keys for further local training. HE serves to conceal each client’s contributions, thereby removing the server’s access to sensitive information. Despite the computational overhead associated with HE, its implementation can markedly enhance patient data security within collaborative learning environments.
While HE mitigates risks such as model inversion or data leakage attributable to compromised servers, it is important to consider that the final models themselves may still retain privacy-sensitive information (see model inversion (Yin et al., 2020) and membership inference (Shokri et al., 2017)). Consequently, integrating additional privacy safeguards, such as differential privacy or partial model sharing, warrants consideration to prevent the potential for memorization of individual training data.
Several FL frameworks implement HE-based solutions for secure federated aggregation (Jin et al., 2023; Cremonesi et al., 2023; Roth et al., 2022), thereby significantly addressing privacy concerns in FL. These frameworks focus on typical applications of horizontal FL and deep learning models. Other approaches include HE in vertical FL applications, such as SecureBoost (Cheng et al., 2021) aimed for financial applications of XGBoost (Chen and Guestrin, 2016) and encrypted Kaplan-Meier for survival analysis in oncology and genome-wide association studies in FL settings (Froelicher et al., 2021).
@@
7.4. Additional defenses against membership inference attacks
Adversarial Regularization (Adv-reg). Nasr et al. (Nasr et al., 2018) proposed a defense that uses similar ideas as GAN. The classifier is trained in conjunction with an MI attacker model. The optimization objective of the target classifier is to reduce the prediction loss while minimizing the MI attack accuracy.
Mixup+MMD. Li et al. (Li et al., 2021) proposed a defense that combines mixup data augmentation and MMD (Maximum Mean Discrepancy (Fortet and Mourier, 1953; Gretton et al., 2012)) based regularization. Instead of training with original instances, mixup data augmentation uses linear combinations of two original instances to train the model. It was shown in (Zhang et al., 2017) that this can improve target model’s generalization. Li et al. (Li et al., 2021) found that they also help to defend against MI attacks. They also proposed to add a regularizer that is the MMD between the loss distribution of members and the loss distribution of a validation set not used in training. This helps make the loss distribution of members to be more similar to the loss distribution on non-members.
Distillation for membership privacy (DMP). Distillation uses labels generated by a teacher model to train a student model. It was proposed in (Hinton et al., 2015) for the purpose of model compression. Shejwalkar et al. (Shejwalkar and Houmansadr, 2021) proposed to use distillation to defend against MI attacks. One first trains a teacher model using the private training set, and then trains a student model using another unlabeled dataset from the same distribution as the private set. The intuition is that since the student model is not directly optimized over the private set, their membership may be protected. the authors also suggested to train a GAN using the private training set and draw samples from the trained GAN to train the student model, when no auxiliary unlabeled data is available.
SELENA. Tang et al. (Tang et al., 2022) proposed a framework named SELENA. One first generates multiple (overlapping) subsets from the training data, then trains one model from each subset. One then generates a new label for each training instance, using the average of predictions generated by models trained without using that instance. Finally, one trains a model using the training dataset using these new labels.
HAMP. Chen et al. (Chen and Pattabiraman, 2023) proposed a defense combining several ideas. First, labels for training instances are made smoother, by changing to and each to , where is a hyperparameter and is the number of classes. Second, an entropy based regularizer is added in the optimization objective. This step is somewhat redundant give that the first step already increases the entropy of the labels. Third, the model does not directly return its output on a queried instance . Instead, one randomly generates another instance, reshuffle the prediction vector of the randomly generated instance based on the order of the probabilities of and returns the reshuffled prediction vector. In essence, this last defense means returning only the order the classes in the prediction vector but not the actual values.
Mem-guard. Jia et al. (Jia et al., 2019) proposed the Mem-guard defense. In this defense, one trains an MI attack model in addition to the target classifier. When the target classifier is queried with an instance, the resulting prediction vector is not directly returned. Instead, one tries to find a perturbed version of the vector such that the perturbation is minimal and does not change the predicted label, and the MI attack model output (0.5,0.5) as its prediction vector.
MIST. Li et al. (Li et al., 2023) proposed the Membership-Invariant Subspace Training (MIST) defense. In this work, the authors showed that subspace learning methods can help defend against membership inference attacks and make the trained model more generalizable. In a typical subspace learning scenario, the whole training set is divided into disjoint subsets and the original model is copied multiple times to generate a submodel set. For each submodel, it is trained using an unique disjoint subset for several optimization steps and then all the submodels are aggregated together to get a new central model. This diversify-aggregate step is repeated until convergence. Based on the subspace learning methods, a new loss called cross-difference loss is proposed. The cross difference loss is defined to be the difference between prediction from the submodel trained with one particular instance and the average predictions from all other submodels trained without this particular instance, since each instance can only be in one subset and thus be used in the training for one submodel. The experiments showed that MIST defense can achieve the state-of-the-art privacy-utility tradeoff. In (Li et al., 2021; Tang et al., 2022; Chen and Pattabiraman, 2023; Li et al., 2023), extensive experiments have shown that several other defenses can provide better empirical privacy-utility tradeoff than DP-SGD.
@@ 8. Looking Ahead: Important Open Questions and Solution Directions
The topic of Federated Learning privacy is seeing a lot of interest in research, development, and deployment. We expect that for the foreseeable future, there will be robust activity on this topic with the balance between research, development, and deployment shifting over time. Here are five high-level directions that will need to be pursued and answered for FL to become truly useful in practice.
@@
Tradeoff between utility and privacy. Most privacy-preserving techniques in FL fall in the tradeoff space between utility of the model (how accurate is the model under benign circumstances) and the privacy afforded to the data. A prototypical example is perturbation of the gradients being reported by the clients. This can be done for example through Differential Privacy (DP). Existing solutions show that the loss in utility is sharp for reasonable levels of privacy protection (Wei et al., 2020; Naseri et al., 2020; Kim et al., 2021). There is work to be done in making this tradeoff more gradual. Broadly, compelling solutions in this space should be able to provide an adjustable tradeoff between these two dimensions.
@@
From centralization to decentralization. The question is will privacy be helped by a move toward greater decentralization, away from the coordination of a central federated server. The direction of decentralized learning has seen some initial results that indicate the answer to the above question is a yes (Koloskova et al., 2020; Sharma et al., 2023; Li et al., 2022b). However, such decentralization also brings in the vulnerability of the learning process to malicious clients. The question of how to ensure security to malicious clients in such learning remains open. The challenge is how to make such a solution, which has some inherent privacy benefits, also secure to misbehaving participants.
Cerification of privacy. We need to be able to certify that the privacy achieved by a given solution meets any given regulation, which holds within a jurisdiction. Such certification is difficult enough in traditional deterministic programs and more so for the stochastic programs that are at the basis of the FL systems. In line with the policy and regulation discussion in Section 5, for FL to become widely adopted in some critical application domains, it would be essential to provide privacy proofs so that the systems can be certified.
Verification of privacy. In one major line of work (Zhao et al., 2023b; Fowl et al., 2022; Pasquini et al., 2022; Boenisch et al., 2023), we have seen that privacy can be violated by the server by sending maliciously crafted models to the clients. This triggers the question what types of malicious models can be examined and detected by the faulty clients. The comparison can be quite a daunting task because of the relative weak computing power of clients, especially mobile and embedded clients, relative to the sizes of the models. The broad question is what does stealthiness of local data exfiltration mean with respect to the capability to verify by the clients.
@@
Regulatory compliance As FL systems become more prevalent, it is imperative to address ethical concerns and ensure transparency in the process towards regulatory compliance of AI systems. This involves understanding the implications of FL on various stakeholders, including data providers, model owners, and end-users. Transparency measures should be implemented to provide visibility into how FL operates, including data usage, model training, and decision-making processes. Additionally, ethical guidelines should be established to govern the responsible development and deployment of FL systems, considering factors such as fairness, bias mitigation, and accountability. Incorporating ethical considerations and transparency measures will enhance trust in FL systems and promote their responsible use in various applications. They are a step towards compliance with emerging AI regulations, such as the EU AI act (Woisetschläger et al., 2024).
@@
Each of these directions presents unique challenges that must be addressed to realize the full potential of FL while safeguarding data privacy.
@@ Acknowledgements.
To all our funding agencies.References
@@
- FLI ([n. d.]) [n. d.]. Federated learning may provide a solution for future digital health challenges, howpublished = https://www.kcl.ac.uk/news/federated-learning-may-provide-a-solution-for-future-digital-health-challenges, note = Accessed: May 15, 2024.
@@
- fet ([n. d.]) [n. d.]. The federated tumor segmentation (fets) initiative. https://www.med.upenn.edu/cbica/fets. Accessed: May 15, 2024.
@@
- hea ([n. d.]) [n. d.]. HealthChain consortium, howpublished = https://www.labelia.org/en/healthchain-project, note = Accessed: May 15, 2024.
@@
- acr ([n. d.]) [n. d.]. Rhino Health and the American College of Radiology Enable Privacy-Preserving Artificial Intelligence. https://www.accesswire.com/728071/rhino-health-and-the-american-college-of-radiology-enable-privacy-preserving-artificial-intelligence. Accessed: May 15, 2024.
@@
- TFD ([n. d.]) [n. d.]. Trustworthy Federated Data Analytics (TFDA). https://tfda.hmsp.center/. Accessed: May 15, 2024.
@@
- Boenisch et al. (2023) Franziska Boenisch, Adam Dziedzic, Roei Schuster, Ali Shahin Shamsabadi, Ilia Shumailov, and Nicolas Papernot. 2023. When the curious abandon honesty: Federated learning is not private. 8th IEEE European Symposium on Security and Privacy (IEEE Euro S&P) (2023).
@@
- Chen and Pattabiraman (2023) Zitao Chen and Karthik Pattabiraman. 2023. Overconfidence is a Dangerous Thing: Mitigating Membership Inference Attacks by Enforcing Less Confident Prediction. arXiv preprint arXiv:2307.01610 (2023).
@@
- Gretton et al. (2012) Arthur Gretton, Karsten M Borgwardt, Malte J Rasch, Bernhard Schölkopf, and Alexander Smola. 2012. A kernel two-sample test. Journal of Machine Learning Research 13, Mar (2012), 723–773.
@@
- Hinton et al. (2015) Geoffrey Hinton, Oriol Vinyals, and Jeff Dean. 2015. Distilling the knowledge in a neural network. arXiv preprint arXiv:1503.02531 (2015).
@@
- Jia et al. (2019) Jinyuan Jia, Ahmed Salem, Michael Backes, Yang Zhang, and Neil Zhenqiang Gong. 2019. MemGuard: Defending against Black-Box Membership Inference Attacks via Adversarial Examples. In Proceedings of the 2019 ACM SIGSAC Conference on Computer and Communications Security. 259–274.
@@
- Li et al. (2023) Jiacheng Li, Ninghui Li, and Bruno Ribeiro. 2023. MIST: Defending Against Membership Inference Attacks Through Membership-Invariant Subspace Training. arXiv preprint arXiv:2311.00919 (2023).
@@
- Luo et al. (2023) B. Luo, Y. Feng, S. Wang, J. Huang, and L. Tassiulas. 2023. Incentive Mechanism Design for Unbiased Federated Learning with Randomized Client Participation. In 2023 IEEE 43rd International Conference on Distributed Computing Systems (ICDCS). IEEE Computer Society, Los Alamitos, CA, USA, 545–555.
@@
- Mansour et al. (2020) Y. Mansour, Mehryar Mohri, Jae Ro, and Ananda Theertha Suresh. 2020. Three Approaches for Personalization with Applications to Federated Learning. ArXiv abs/2002.10619 (2020). https://api.semanticscholar.org/CorpusID:211296702
@@
- Nasr et al. (2018) Milad Nasr, Reza Shokri, and Amir Houmansadr. 2018. Machine learning with membership privacy using adversarial regularization. In Proceedings of the 2018 ACM SIGSAC Conference on Computer and Communications Security. ACM, 634–646.
@@
- Roth and Lempel (1989) Ron M Roth and Abraham Lempel. 1989. On MDS codes via Cauchy matrices. IEEE transactions on information theory 35, 6 (1989), 1314–1319.
@@
- Sharma et al. (2023) Atul Sharma, Joshua C Zhao, Wei Chen, Qiang Qiu, Saurabh Bagchi, and Somali Chaterji. 2023. How to Learn Collaboratively-Federated Learning to Peer-to-Peer Learning and What’s at Stake. In 2023 53rd Annual IEEE/IFIP International Conference on Dependable Systems and Networks-Supplemental Volume (DSN-S). IEEE, 122–126.
- Shejwalkar and Houmansadr (2021) Virat Shejwalkar and Amir Houmansadr. 2021. Membership privacy for machine learning models through knowledge transfer. In Proceedings of the AAAI Conference on Artificial Intelligence.
@@
- Tang et al. (2021) Tingting Tang, Ramy E Ali, Hanieh Hashemi, Tynan Gangwani, Salman Avestimehr, and Murali Annavaram. 2021. Verifiable coded computing: Towards fast, secure and private distributed machine learning. arXiv preprint arXiv:2107.12958 (2021).
- Tang et al. (2022) Xinyu Tang, Saeed Mahloujifar, Liwei Song, Virat Shejwalkar, Milad Nasr, Amir Houmansadr, and Prateek Mittal. 2022. Mitigating membership inference attacks by Self-Distillation through a novel ensemble architecture. In 31st USENIX Security Symposium (USENIX Security 22). 1433–1450.
@@
- Yin et al. (2020) Hongxu Yin, Pavlo Molchanov, Jose M Alvarez, Zhizhong Li, Arun Mallya, Derek Hoiem, Niraj K Jha, and Jan Kautz. 2020. Dreaming to distill: Data-free knowledge transfer via deepinversion. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8715–8724.
@@
- Zhang et al. (2017) Hongyi Zhang, Moustapha Cisse, Yann N Dauphin, and David Lopez-Paz. 2017. mixup: Beyond empirical risk minimization. arXiv preprint arXiv:1710.09412 (2017).
@@
- Zhang et al. (2021) Tuo Zhang, Chaoyang He, Tianhao Ma, Lei Gao, Mark Ma, and Salman Avestimehr. 2021. Federated Learning for Internet of Things. In Proceedings of the 19th ACM Conference on Embedded Networked Sensor Systems (Coimbra, Portugal) (SenSys ’21). Association for Computing Machinery, New York, NY, USA, 413–419. https://doi.org/10.1145/3485730.3493444
@@
- Zhao et al. (2024) Joshua Christian Zhao, Ahaan Dabholkar, Atul Sharma, and Saurabh Bagchi. 2024. Leak and Learn: An Attacker’s Cookbook to Train Using Leaked Data from Federated Learning. In Conference on Computer Vision and Pattern Recognition (CVPR). IEEE Computer Society, 1–10.

### v2

@@ 7.3. Homomorphic Encryption
Clients can employ HE to encrypt their model updates within the FL paradigm before transmission. Consequently, the server exclusively receives encrypted updates and is precluded from accessing any raw model updates. Utilizing an HE protocol, the server aggregates only encrypted weights, preserving privacy. Following aggregation, the still-encrypted updated global model is sent back to the clients. Subsequently, the clients decrypt the model using their keys for further local training. HE serves to conceal each client’s contributions, thereby removing the server’s access to sensitive information. Despite the computational overhead associated with HE, its implementation can markedly enhance patient data security within collaborative learning environments.
Although HE mitigates risks such as model inversion or data leakage attributable to compromised servers, it is important to consider that the final models themselves may still retain privacy-sensitive information (see membership inference in Section 6.2). Consequently, integrating additional privacy safeguards, such as differential privacy or partial model sharing, should be considered to prevent the potential for memorization of individual training data.
Several FL frameworks implement HE-based solutions for secure federated aggregation (Jin et al., 2023; Cremonesi et al., 2023; Roth et al., 2022), thereby significantly improving privacy concerns in FL. These frameworks focus on typical applications of horizontal FL and deep learning models. Other approaches include HE in vertical FL applications, such as SecureBoost (Cheng et al., 2021), which aims for financial applications of XGBoost (Chen and Guestrin, 2016) and encrypted Kaplan-Meier for survival analysis in oncology and genome-wide association studies in FL settings (Froelicher et al., 2021).
@@
7.4. Discussions
Measurements for success. Unlike attacks, it is quite difficult to directly compare the different privacy-preserving defenses in FL with each other. For example, while all data reconstruction attacks ultimately aim to reconstruct or recover private user data, the goals of each defense in FL are inherently different. For example, secure aggregation only aims at preventing participants (including the server) from seeing individual updates while providing no guarantees beyond this. Differential privacy, on the other hand, does not encrypt individual updates but instead provides theoretical guarantees on the effectiveness of data reconstruction or membership inference attacks. In this way, there is no universal metric that can be used to directly compare the different defenses with each other. Instead, the defenses should simply be evaluated on what they are claiming to defend against(e.g., a secure aggregation method would be a failure if an attacker could simply unencrypt individual updates). When choosing different defense methods for deployment, designers should also keep in mind the differing objectives of each defense.
Maintaining success of defenses in real-world deployment is difficult. Success in deployment would be the case where user data privacy is never breached, as any breach of privacy would be a critical problem. This is especially true given the large-scale nature of cross-device FL or the extremely sensitive nature of cross-silo FL deployments (i.e., hospital or financial data). Given how crucial it is to prevent privacy problems, it is often better to rely on certifiable defenses. An example would be differential privacy, which is currently being used for Google Gboard (Xu et al., 2023).
Powerful attackers. Recent research has shown that more powerful privacy attacks are possible in FL. In this case, individual defenses may not be enough. For example, the guarantee of secure aggregation is not enough against an attacker who can scale regardless of the number of clients (Zhao et al., 2023b) and differential privacy alone is not enough to stop an attacker from learning a representation of the data (Hitaj et al., 2017). Given knowledge of adversary capabilities, it could even be possible to decrease the efficiency of certain data reconstruction attacks by exploiting data heterogeneity (Zhao et al., 2023b) (current defenses work regardless of what is happening with client data heterogeneity). A more effective method against these attacks could be the use of both secure aggregation and differential privacy. However, even this combination may not be enough. If malicious clients can form a large majority of participating clients, it is still possible to single out individual updates and leak information (Boenisch et al., 2023a). Preventing these additional attacks will require more careful client selection or even stronger differential privacy guarantees, which can often come at the cost of model performance.
@@ 8. Looking Ahead: Important Open Questions and Solution Directions
Compared to traditional machine learning, Federated Learning (FL) provides a unique perspective on security and privacy issues. Although learning seems to be more private due to the data being held locally on the clients, at the same time, the process still opens up new avenues for attacks due to the sharing of machine learning updates. Given inherent restrictions of FL, such as client computation and communication capabilities or other drawbacks such as in model performance when using certain defenses, research will need to continue in privacy-preserving FL. As a privacy-preserving technology by design, it is essential for FL to truly preserve the privacy of user data, even more than traditional (i.e., centralized) machine learning. At the same time, given the importance of model accuracy and performance in practice, it is also essential for FL to balance the trade-off between model utility and privacy.
@@
The topic of Federated Learning privacy is attracting a lot of interest in research, development, and deployment. We expect that for the foreseeable future there will be robust activity on this topic with the balance between research, development, and deployment shifting over time. Here are six high-level directions that will need to be pursued and answered for FL to become truly useful in practice.
@@
Tradeoff between utility and privacy. Most privacy-preserving techniques in FL fall in the tradeoff space between utility of the model (how accurate is the model under benign circumstances) and the privacy afforded to the data. A prototypical example is perturbation of the gradients being reported by the clients. This can be done through differential privacy (DP). Existing solutions show that the loss in utility is sharp for reasonable levels of privacy protection (Wei et al., 2020; Naseri et al., 2020; Kim et al., 2021). There is work to be done in making this tradeoff more gradual. Broadly, compelling solutions in this space should be able to provide an adjustable tradeoff between these two dimensions.
@@
From centralization to decentralization. The big open question is will privacy be helped by a move toward greater decentralization, away from the coordination of a central federated server. The direction of decentralized learning has seen some initial results indicating a promising answer (Koloskova et al., 2020; Sharma et al., 2023b; Li et al., 2022b). However, such decentralization also brings in the vulnerability of the learning process to malicious clients (Sharma et al., 2023a). The question of how to ensure security to malicious clients in such learning remains open. The challenge is how to make a solution that has inherent privacy benefits also secure to misbehaving participants.
@@
Regulatory and ethical compliance. We need to be able to certify that the privacy achieved by a given solution meets a given regulation, which applies within a jurisdiction. In line with the policy and regulation discussion in Section 5, for FL to become widely adopted in some critical application domains, it would be essential to provide privacy proofs so that the systems can be certified. Such certification is difficult enough in traditional deterministic programs and more so for the stochastic programs that are the basis of FL systems. As FL systems become more prevalent, it is also imperative to address ethical concerns and ensure transparency in the process towards regulatory compliance of AI systems. This involves understanding the implications of FL on various stakeholders, including data providers, model owners, and end users. Transparency measures should be implemented to provide visibility into how FL operates, including data usage, model training, and decision-making processes. Additionally, ethical guidelines should be established to govern the responsible development and deployment of FL systems, considering factors such as fairness, bias mitigation, and accountability. Incorporating ethical considerations and transparency measures will enhance trust in FL systems and promote their responsible use in various applications. They are a step towards compliance with emerging AI regulations, such as the EU AI act (Woisetschläger et al., 2024).
Verification of privacy by clients. In one major line of work (Zhao et al., 2023b; Fowl et al., 2022; Pasquini et al., 2022; Boenisch et al., 2023b), we have seen that privacy can be violated by the server by sending maliciously crafted models to the clients. This triggers the question of how malicious models can be examined and detected by the clients. The comparison can be quite a daunting task because of the relative weak computing power of clients, especially mobile and embedded clients, relative to the sizes of the models. The broad question is what the stealthiness of data exfiltration attacks implies for the capability to verify manipulated models by the clients.
Handling diverse data sources and nodes. In FL, the data sources are often diverse in various aspects, such as kinds of data (images of different resolutions, videos of different lengths, etc.) and level of noise in the data (naturally present or adversarially introduced). It is imperative for private FL to be able to work with such diverse data sources, in line with the huge development of multimodal AI models. Another dimension of diversity comes from the nodes themselves, which can differ in computational power and communication connectivity. It is an ongoing challenge to enable FL with such diverse nodes, ensuring that FL does not bias the model in favor of well-connected nodes and thus leak more information from them.
Scalability of privacy-preserving decentralized learning. A prime motivation for moving toward decentralized learning is that solutions are more scalable (to a larger number of nodes and larger models) than in traditional centralized learning. However, when one interposes privacy-preserving techniques, this scalability becomes more challenging to maintain. Consider that several privacy-enhancing techniques, such as differential privacy, secure aggregation, and secure multi-party computation, which are prime tools in private FL, can introduce significant computation or communication overheads. Thus, it is an open question how ongoing advances in efficiency of these privacy-preserving techniques can be incorporated into private FL.
Personalized FL and nonIID data. A further challenge arises from the fact that client data in FL is typically nonIID (i.e., not identically distributed across clients), which can slow or hinder convergence in standard FL approaches. Personalized FL tackles this by allowing each client to maintain a set of local parameters reflecting its unique data distribution while still benefiting from a global shared model. This technique can significantly improve performance for clients whose data differ markedly from the overall population. However, ensuring privacy under personalization remains an open research area, especially when balancing the need for local customization with the constraints of private communication and aggregation. The tension between personalization, privacy, and efficient convergence continues to drive new methods in this space.
@@
Each of these directions presents unique challenges that must be addressed to realize the full potential of FL while safeguarding data privacy. We look forward to a field with research advances working hand-in-hand with robust practical implementations and strong policy, all leading to widespread use of Federated Learning in security and privacy-critical applications.
@@ Acknowledgements.
This material is based in part upon work supported by: (i) S. Bagchi and J. Zhao: Cisco Research, the DEVCOM ARL Army Research Office under Contract number W911NF-2020-221, and the National Science Foundation under Grant Numbers CNS-2333487 and CNS-2038986. (ii) J. Li and N. Li: Cisco Research and the National Science Foundation under Grant Numbers CNS-2247794 and CNS-2207204. (iii) S. Chaterji: DEVCOM ARL Army Research Office under Contract number W911NF-2020-221 and NSF Grant Numbers CNS-2146449 (NSF-CAREER) and CNS-2333487 (NSF-FRONTIER). The first author Joshua Zhao was supported through part of his PhD by the Andrews Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.References
@@
- FLI ([n. d.]) [n. d.]. Federated learning may provide a solution for future digital health challenges, howpublished = https://www.kcl.ac.uk/news/federated-learning-may-provide-a-solution-for-future-digital-health-challenges, note = Accessed: March 13, 2025.
@@
- fet ([n. d.]) [n. d.]. The federated tumor segmentation (fets) initiative. https://www.med.upenn.edu/cbica/fets. Accessed: March 13, 2025.
@@
- hea ([n. d.]) [n. d.]. HealthChain consortium, howpublished = https://www.labelia.org/en/healthchain-project, note = Accessed: March 13, 2025.
@@
- acr ([n. d.]) [n. d.]. Rhino Health and the American College of Radiology Enable Privacy-Preserving Artificial Intelligence. https://www.accesswire.com/728071/rhino-health-and-the-american-college-of-radiology-enable-privacy-preserving-artificial-intelligence. Accessed: March 13, 2025.
@@
- TFD ([n. d.]) [n. d.]. Trustworthy Federated Data Analytics (TFDA). https://tfda.hmsp.center/. Accessed: March 13, 2025.
@@
- Imp (2021) 2021. Application of Federated Learning in Predictive Maintenance to Predict Remaining Useful Life. (2021). https://ifoadatascienceresearch.github.io/assets/pdfs/FL_predictive_maintenance.pdf
@@
- Augenstein et al. (2022) Sean Augenstein, Andrew Hard, Lin Ning, Karan Singhal, Satyen Kale, Kurt Partridge, and Rajiv Mathews. 2022. Mixed Federated Learning: Joint Decentralized and Centralized Learning. arXiv:2205.13655
@@
- Boenisch et al. (2023a) Franziska Boenisch, Adam Dziedzic, Roei Schuster, Ali Shahin Shamsabadi, Ilia Shumailov, and Nicolas Papernot. 2023a. Reconstructing individual data points in federated learning hardened with differential privacy and secure aggregation. In 2023 IEEE 8th European Symposium on Security and Privacy (EuroS&P). IEEE, 241–257.
- Boenisch et al. (2023b) Franziska Boenisch, Adam Dziedzic, Roei Schuster, Ali Shahin Shamsabadi, Ilia Shumailov, and Nicolas Papernot. 2023b. When the curious abandon honesty: Federated learning is not private. 8th IEEE European Symposium on Security and Privacy (IEEE Euro S&P) (2023).
@@
- Chaterji et al. (2019) Somali Chaterji, Jinkyu Koo, Ninghui Li, Folker Meyer, Ananth Grama, and Saurabh Bagchi. 2019. Federation in genomics pipelines: techniques and challenges. Briefings in bioinformatics 20, 1 (2019), 235–244.
@@
- Cho et al. (2023) Yae Jee Cho, Gauri Joshi, and Dimitrios Dimitriadis. 2023. Local or Global: Selective Knowledge Assimilation for Federated Learning with Limited Labels. In 2023 IEEE/CVF International Conference on Computer Vision (ICCV). 17041–17050. https://doi.org/10.1109/ICCV51070.2023.01567
@@
- De et al. (2022) Soham De, Leonard Berrada, Jamie Hayes, Samuel L. Smith, and Borja Balle. 2022. Unlocking High-Accuracy Differentially Private Image Classification through Scale. CoRR abs/2204.13650 (2022).
@@
- Ge et al. (2021) Ning Ge, Guanghao Li, Li Zhang, and Yi Liu. 2021. Failure prediction in production line based on federated learning: an empirical study. Journal of Intelligent Manufacturing 33 (2021), 2277 – 2294. https://api.semanticscholar.org/CorpusID:231719718
@@
- Jain et al. (2021) Pranjal Jain, Shreyas Goenka, Saurabh Bagchi, Biplab Banerjee, and Somali Chaterji. 2021. Federated action recognition on heterogeneous embedded devices. arXiv preprint arXiv:2107.12147 (2021).
@@
- Mahgoub et al. (2019) Ashraf Mahgoub, Paul Wood, Alexander Medoff, Subrata Mitra, Folker Meyer, Somali Chaterji, and Saurabh Bagchi. 2019. SOPHIA: Online reconfiguration of clustered NoSQL databases for Time-Varying workloads. In 2019 USENIX Annual Technical Conference (USENIX ATC 19). 223–240.
@@
- Sander et al. (2023) Tom Sander, Pierre Stock, and Alexandre Sablayrolles. 2023. TAN Without a Burn: Scaling Laws of DP-SGD. In International Conference on Machine Learning, ICML 2023, 23-29 July 2023, Honolulu, Hawaii, USA (Proceedings of Machine Learning Research, Vol. 202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.). PMLR, 29937–29949.
@@
- Sharma et al. (2023a) Atul Sharma, Wei Chen, Joshua Zhao, Qiang Qiu, Saurabh Bagchi, and Somali Chaterji. 2023a. Flair: Defense against model poisoning attack in federated learning. In Proceedings of the 2023 ACM Asia Conference on Computer and Communications Security. 553–566.
- Sharma et al. (2021) Atul Sharma, Wei Chen, Joshua Zhao, Qiang Qiu, Somali Chaterji, and Saurabh Bagchi. 2021. TESSERACT: Gradient Flip Score to Secure Federated Learning Against Model Poisoning Attacks. arXiv:2110.10108 [cs.LG] https://arxiv.org/abs/2110.10108
- Sharma et al. (2023b) Atul Sharma, Joshua C Zhao, Wei Chen, Qiang Qiu, Saurabh Bagchi, and Somali Chaterji. 2023b. How to Learn Collaboratively-Federated Learning to Peer-to-Peer Learning and What’s at Stake. In 2023 53rd Annual IEEE/IFIP International Conference on Dependable Systems and Networks-Supplemental Volume (DSN-S). IEEE, 122–126.
@@
- Sun et al. (2024) Hao Sun, Xiaoli Tang, Chengyi Yang, Zhenpeng Yu, Xiuli Wang, Qijie Ding, Zengxiang Li, and Han Yu. 2024. HiFi-Gas: Hierarchical Federated Learning Incentive Mechanism Enhanced Gas Usage Estimation. Proceedings of the AAAI Conference on Artificial Intelligence 38, 21 (Mar. 2024), 22824–22832. https://ojs.aaai.org/index.php/AAAI/article/view/30317
@@
- Wang et al. (2021) Zhiguo Wang, Xintong Wang, Ruoyu Sun, and Tsung-Hui Chang. 2021. Federated Semi-Supervised Learning with Class Distribution Mismatch. CoRR abs/2111.00010 (2021).
@@
- Xu et al. (2023) Zheng Xu, Yanxiang Zhang, Galen Andrew, Christopher A Choquette-Choo, Peter Kairouz, H Brendan McMahan, Jesse Rosenstock, and Yuanbo Zhang. 2023. Federated learning of gboard language models with differential privacy. arXiv preprint arXiv:2305.18465 (2023).
@@
- Zhang et al. (2021a) Hongyi Zhang, Jan Bosch, and Helena Holmström Olsson. 2021a. Real-time End-to-End Federated Learning: An Automotive Case Study. In 2021 IEEE 45th Annual Computers, Software, and Applications Conference (COMPSAC). 459–468. https://doi.org/10.1109/COMPSAC51774.2021.00070
@@
- Zhang et al. (2021b) Tuo Zhang, Chaoyang He, Tianhao Ma, Lei Gao, Mark Ma, and Salman Avestimehr. 2021b. Federated Learning for Internet of Things. In Proceedings of the 19th ACM Conference on Embedded Networked Sensor Systems (Coimbra, Portugal) (SenSys ’21). Association for Computing Machinery, New York, NY, USA, 413–419. https://doi.org/10.1145/3485730.3493444
@@
- Zhao et al. (2024) Joshua C Zhao, Ahaan Dabholkar, Atul Sharma, and Saurabh Bagchi. 2024. Leak and Learn: An Attacker’s Cookbook to Train Using Leaked Data from Federated Learning. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 12247–12256.